# The asymptotic Plancherel formula and Lusztig’s asymptotic algebra for $\tilde{\mathsf{A}}_{n}$

### v1

@@ 5 The asymptotic Plancherel Theorem
In this section we prove an asymptotic Plancherel Theorem for type $\tilde{\mathsf{A}}_{n}$. This notion was introduced in [11, 12], and Proposition 5.4 confirms [11, Conjecture 9.8] and [10, Conjecture 5.15] for affine type $\mathsf{A}$. The asymptotic Plancherel Theorem will play a crucial role, often leading to efficient proofs of certain statements (for example, see Theorems 5.7, 5.8, and Claim 3 in the proof of Theorem 7.6). We note that other interesting applications of the Plancherel Theorem in Kazhdan-Lusztig theory have recently been made by Dawydiak [5].
@@ 5.1 The Plancherel Theorem for $\tilde{\mathsf{A}}_{n}$
We first recall the Plancherel Theorem (from [2, 24]). The canonical trace on $\widetilde{H}$ is the linear function $\operatorname{\mathsf{Tr}}:\widetilde{H}\to\mathsf{R}$ with $\operatorname{\mathsf{Tr}}(T_{w})=\delta_{w,e}$. Induction on $\ell(v)$ shows that $\operatorname{\mathsf{Tr}}(T_{u}T_{v^{-1}})=\delta_{u,v}$ for all $u,v\in\widetilde{W}$. This implies that $\operatorname{\mathsf{Tr}}(h_{1}h_{2})=\operatorname{\mathsf{Tr}}(h_{2}h_{1})$ for $h_{1},h_{2}\in\widetilde{H}$, and that
$\langle h_{1},h_{2}\rangle=\operatorname{\mathsf{Tr}}(h_{1}h_{2}^{*})$
defines an inner product on $\widetilde{H}$, with $\langle T_{u},T_{v}\rangle=\delta_{u,v}$ (recall here that $\ast$ is defined in (4.2)).
The Plancherel Theorem is a spectral decomposition of the canonical trace functional (see [24]). We now explicitly describe this decomposition in type $\tilde{\mathsf{A}}_{n}$. Recall that $\mathsf{v}=-\mathsf{q}^{-1}$. Let
@@
The Plancherel Theorem is usually expressed as an analytic statement, and some comments are required for the translation between Theorem 5.1 and this analytic statement from [2, 24]. Specialise $\mathsf{q}\to q$ with $q$ a real number with $q>1$, and extend scalars of the Hecke algebra to $\mathbb{C}$. Let $\mathbb{T}$ be the group of complex numbers of modulus $1$. For $\lambda\vdash n+1$ let $dz_{\lambda}$ be normalised Haar measure on the group
$\mathbb{T}_{\lambda}=\{(z_{1},z_{2},\ldots,z_{r(\lambda)})\in\mathbb{T}^{r(\lambda)}\mid z_{1}^{\lambda_{1}}z_{2}^{\lambda_{2}}\cdots z_{r(\lambda)}^{\lambda_{r(\lambda)}}=1\}.$
On specialising the “variables” $z_{1},\ldots,z_{r(\lambda)}$ in our representations to complex numbers of modulus $1$ with $z_{1}^{\lambda_{1}}z_{2}^{\lambda_{2}}\cdots z_{r(\lambda)}^{\lambda_{r(\lambda)}}=1$, properties of the Haar measure give
@@
and $c^{(3)}(\zeta_{(3)})=1$. We have $|G_{(1,1,1)}|=6$ and $|G_{(2,1)}|=|G_{(3)}|=1$. For the analytic formulation of the Plancherel Theorem, note that if $\lambda=(3)$ then $\mathbb{T}_{\lambda}=\{z\in\mathbb{C}\mid z^{3}=1\}=\{1,\omega,\omega^{2}\}$ where $\omega=e^{2\pi i/3}$, and so the normalised Haar measure on $\mathbb{T}_{\lambda}$ is the discrete measure assigning mass $1/3$ to each atom, and we recover the analytic formulation of the Plancherel Theorem for $\tilde{\mathsf{A}}_{2}$ computed explicitly in [25, Theorem 4.3].
@@ 5.2 Asymptotic Plancherel Theorem
In this section we prove an asymptotic version of the Plancherel Theorem for type $\tilde{\mathsf{A}}_{n}$, confirming [11, Conjecture 9.8] and [10, Conjecture 5.15] for affine type $\mathsf{A}$. The existence of such a formula relies on a certain compatibility between the degree (in $\mathsf{q}$) of the Plancherel measure and the bound of the representations $\pi_{\lambda}$, proved in Proposition 5.4 and Corollary 5.5. Recall the definition of the specialisation $\mathrm{sp}_{0}$ from Section 4.3.
@@
Recall the definition of the leading matrices, the ring $\mathfrak{C}_{\lambda}$ from Section 4.3 and the $\ast$-operation on $\mathfrak{C}_{\lambda}$ in Section 4.2. By Lemmas 4.6 and 4.11 we have
$\mathfrak{c}_{\lambda}(w^{-1})=\mathfrak{c}_{\lambda}(w)^{*}.$ (5.2)
We call the following theorem the Asymptotic Plancherel Theorem.
Theorem 5.6.
For $A,B\in\mathfrak{C}_{\lambda}$ let
$\langle A,B\rangle_{\lambda}^{\infty}=\frac{1}{|G_{\lambda}|}\bigg{[}\operatorname{\mathsf{tr}}(AB^{*})\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})\bigg{]}_{\mathrm{ct}}.$
Then $\langle\cdot,\cdot\rangle_{\lambda}^{\infty}$ is an inner product on the $\mathbb{Z}$-module $\mathfrak{C}_{\lambda}$, and the elements $\mathfrak{c}_{\lambda}(w)$, $w\in\Delta_{\lambda}$, form an orthonormal basis. Moreover, $\langle AB,C\rangle_{\lambda}^{\infty}=\langle B,A^{*}C\rangle_{\lambda}^{\infty}$.
Proof.
It is clear that $\langle\cdot,\cdot\rangle_{\lambda}^{\infty}$ is bilinear. Let $A^{t}$ denote matrix transpose. Since
$\operatorname{\mathsf{tr}}(BA^{*})=\operatorname{\mathsf{tr}}((BA^{*})^{t})=\operatorname{\mathsf{tr}}((A^{*})^{t}B^{t})=\mathrm{conj}(\operatorname{\mathsf{tr}}(AB^{*}))$
and $\mathrm{conj}\big{(}\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})\big{)}=\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})$ the form $\langle\cdot,\cdot\rangle_{\lambda}^{\infty}$ is symmetric.
Let $u,v\in\Delta_{\lambda}$. Since $\langle T_{u},T_{v}\rangle=\delta_{u,v}$ the Plancherel Theorem (Theorem 5.1) gives
$\delta_{u,v}=\sum_{\mu\vdash n+1}\langle T_{u},T_{v}\rangle_{\mu},$
and by Corollary 5.5 each specialisation $\mathrm{sp}_{0}\big{(}\langle T_{u},T_{v}\rangle_{\mu}\big{)}$ exists, and is zero unless $\mu=\lambda$. Thus $\delta_{u,v}=\mathrm{sp}_{0}\big{(}\langle T_{u},T_{v}\rangle_{\lambda}\big{)}.$ Now, by Theorem 4.7 we have
$\chi_{\lambda}(T_{u}T_{v^{-1}})=\mathsf{q}^{2\ell(\mathsf{w}_{\lambda^{\prime}})}\operatorname{\mathsf{tr}}(\mathfrak{c}_{\lambda}(u)\mathfrak{c}_{\lambda}(v)^{*})+(\textrm{terms of strictly lower degree}),$
@@
$C_{\lambda}(\mathsf{q})\frac{\chi_{\lambda}(T_{u}T_{v^{-1}})}{c^{\lambda}(\zeta_{\lambda})c^{\lambda}(\zeta_{\lambda}^{-1})}$ $=\operatorname{\mathsf{tr}}(\mathfrak{c}_{\lambda}(u)\mathfrak{c}_{\lambda}(v)^{*})\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})+(\textrm{terms of degree $<0$}).$
@@
$\delta_{u,v}=\mathrm{sp}_{0}\big{(}\langle T_{u},T_{v}\rangle_{\lambda}\big{)}=\frac{1}{|G_{\lambda}|}\bigg{[}\operatorname{\mathsf{tr}}(\mathfrak{c}_{\lambda}(u)\mathfrak{c}_{\lambda}(v)^{*})\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})\bigg{]}_{\mathrm{ct}}=\langle\mathfrak{c}_{\lambda}(u),\mathfrak{c}_{\lambda}(v)\rangle_{\lambda}^{\infty}.$
If $A=\sum_{w\in\Delta_{\lambda}}a_{w}\mathfrak{c}_{\lambda}(w)$ with $A\neq 0$ it follows that $\langle A,A\rangle_{\lambda}^{\infty}=\sum_{w\in\Delta_{\lambda}}a_{w}^{2}>0$, and so $\langle\cdot,\cdot\rangle_{\lambda}^{\infty}$ is positive definite, and the elements $\mathfrak{c}_{\lambda}(w)$, $w\in\Delta_{\lambda}$, are orthonormal. Finally, since $\operatorname{\mathsf{tr}}((AB)C^{*})=\operatorname{\mathsf{tr}}(BC^{*}A)=\operatorname{\mathsf{tr}}(B(A^{*}C)^{*})$ we have $\langle AB,C\rangle_{\lambda}^{\infty}=\langle B,A^{*}C\rangle_{\lambda}^{\infty}$. ∎
@@
By Corollary 4.8, it only remains to show that if $w\in\Delta_{\lambda}$ then $\deg\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})=\ell(\mathsf{w}_{\lambda^{\prime}})$. If $w\in\Delta_{\lambda}$ then by the Asymptotic Plancherel Theorem (Theorem 5.6) we have $1=\langle\mathfrak{c}_{\lambda}(w),\mathfrak{c}_{\lambda}(w)\rangle_{\lambda}^{\infty}$ and thus $\deg\langle T_{w},T_{w}\rangle_{\lambda}=0$, and the result follows from Corollary 5.5. ∎
@@
By Proposition 4.12 the map $\psi:\mathcal{J}_{\lambda}\to\mathfrak{C}_{\lambda},\mathsf{t}_{w}\mapsto\mathfrak{c}_{\lambda}(w)$ is a surjective ring homomorphism. To prove that $\psi$ is injective, if $a=\sum_{w\in\Delta_{\lambda}}a_{w}\mathsf{t}_{w}\in\mathcal{J}_{\lambda}$ with $\psi(a)=0$ then for each $v\in\Delta_{\lambda}$ we have
$\sum_{w\in\Delta_{\lambda}}a_{w}\mathfrak{c}_{\lambda}(w)\mathfrak{c}_{\lambda}(v)^{*}=0.$
Taking traces, multiplying by $\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})$, and applying Theorem 5.6 gives $a_{v}=0$. ∎
@@
for $A,B\in\mathfrak{C}_{\lambda}$. We call $d\mu_{\lambda}^{\infty}(z_{\lambda})$ the asymptotic Plancherel measure.

### v2

@@ 4.3 Leading matrices and the characters $\chi_{\lambda}^{\infty}$ of $\mathcal{J}$
It follow from Lemmas 4.6 and 4.11 that
$\mathfrak{c}_{\lambda}(w^{-1})=\mathfrak{c}_{\lambda}(w)^{*}$ (4.5)
where, as in Section 4.3, $\mathfrak{c}_{\lambda}(w)^{*}$ denotes the conjugate transpose matrix.
For $\lambda\vdash n+1$ define $\pi_{\lambda}^{\infty}:\mathcal{J}\to\mathrm{Mat}_{N_{\lambda}}(\mathbb{Z}[\zeta_{\lambda}])$ by
$\pi_{\lambda}^{\infty}(A)=\sum_{w\in\Delta_{\lambda}}a_{w}\mathfrak{c}_{\lambda}(w)\quad\text{where}\quad A=\sum_{w\in\widetilde{W}}a_{w}\mathsf{t}_{w}.$ (4.6)
In particular, note that $\pi_{\lambda}^{\infty}(\mathsf{t}_{w})=0$ if $w\notin\Delta_{\lambda}$. Proposition 4.12 shows that $\pi_{\lambda}^{\infty}$ gives a matrix representation of $\mathcal{J}$. Let $\chi_{\lambda}^{\infty}$ be the character of $\pi_{\lambda}^{\infty}$, and so
$\chi_{\lambda}^{\infty}(A)=\operatorname{\mathsf{tr}}(\pi_{\lambda}^{\infty}(A))\quad\text{for $A\in\mathcal{J}$}$
where $\operatorname{\mathsf{tr}}$ denotes matrix trace.
@@ 5 The asymptotic Plancherel Theorem
In this section we prove an asymptotic Plancherel Theorem for type $\tilde{\mathsf{A}}_{n}$. This notion was introduced in [14, 15], and Proposition 5.8 confirms [14, Conjecture 9.8] and [13, Conjecture 5.15] for affine type $\mathsf{A}$. The asymptotic Plancherel Theorem will play a crucial role, often leading to efficient proofs of certain statements (for example, see Theorems 5.10, 5.11, and Claim 3 in the proof of Theorem 7.6). We note that other interesting applications of the Plancherel Theorem in Kazhdan-Lusztig theory have recently been made by Dawydiak [6].
5.1 The canonical trace on $\mathcal{J}$
In this subsection (and only in this subsection) let $(W,S)$ be a Coxeter system of arbitrary type, let $\varphi:W\to\mathbb{Z}_{\geq 0}$ be a weight function with $\varphi(s)>0$, and let $H$ be the associated weighted Hecke algebra over $\mathsf{R}=\mathbb{Z}[\mathsf{q},\mathsf{q}^{-1}]$, as in [5, 24]. We assume that the weighted Coxeter system $(W,S,\varphi)$ is bounded as in [24, §13.2], and that properties P1–P15 hold. Thus the asymptotic algebra $\mathcal{J}$ of $H$ is an associative $\mathbb{Z}$-algebra (see [24, §18.3]).
Recall the definition of the set $\mathcal{D}$ of distinguished involutions, and the integers $n_{d}$ $(d\in\mathcal{D}$) from Section 1.6. For $w\in W$ we write $d(w)$ for the unique element of $\mathcal{D}$ in the right cell containing $w$ (c.f. P13). By P5 we have $n_{d(x)}=\gamma_{x,x^{-1},d(x)}=\pm 1.$
Recall that the canonical trace on $H$ is the $\mathsf{R}$-linear function $\operatorname{\mathsf{Tr}}:H\to\mathsf{R}$ with $\operatorname{\mathsf{Tr}}(T_{w})=\delta_{w,e}$ (see [11, §8.1]). Induction on $\ell(v)$ shows that $\operatorname{\mathsf{Tr}}(T_{u}T_{v^{-1}})=\delta_{u,v}$ for all $u,v\in W$. This implies that $\operatorname{\mathsf{Tr}}(h_{1}h_{2})=\operatorname{\mathsf{Tr}}(h_{2}h_{1})$ for $h_{1},h_{2}\in H$, and that
$\langle h_{1},h_{2}\rangle=\operatorname{\mathsf{Tr}}(h_{1}h_{2}^{*})$
defines a nondegenerate symmetric bilinear form on $H$, with $\langle T_{u},T_{v}\rangle=\delta_{u,v}$ (recall here that $\ast$ is defined in (4.2)). We now define a canonical trace on the asymptotic algebra $\mathcal{J}$, following the finite case given in [24, §20.1(b)].
Definition 5.1.
Define a linear map $\operatorname{\mathsf{Tr}}^{\infty}:\mathcal{J}\to\mathbb{Z}$ and a bilinear form $\langle\cdot,\cdot\rangle^{\infty}:\mathcal{J}\times\mathcal{J}\to\mathbb{Z}$ by
$\operatorname{\mathsf{Tr}}^{\infty}\left(\sum a_{w}\mathsf{t}_{w}\right)=\sum_{d\in\mathcal{D}}n_{d}a_{d}\quad\text{and}\quad\langle A,B\rangle^{\infty}=\operatorname{\mathsf{Tr}}^{\infty}(AB^{*})$
for $A,B\in\mathcal{J}$, where $\left(\sum a_{w}\mathsf{t}_{w}\right)^{*}=\sum a_{w}\mathsf{t}_{w^{-1}}$.
Theorem 5.2.
With the above notation and assumptions, we have:
  1. $(1)$
    The linear map $\operatorname{\mathsf{Tr}}^{\infty}$ is a trace functional on $\mathcal{J}$.
  2. $(2)$
    The bilinear form $\langle\cdot,\cdot\rangle^{\infty}$ is an inner product and $(\mathsf{t}_{w})_{w\in W}$ is an orthonormal basis of $\mathcal{J}$.
  3. $(3)$
    For $A,B,C\in\mathcal{J}$ we have $\langle AB,C\rangle^{\infty}=\langle B,A^{*}C\rangle^{\infty}$.
Proof.
Since $\mathsf{t}_{x}\mathsf{t}_{y}^{*}=\sum_{z\in W}\gamma_{x,y^{-1},z^{-1}}\mathsf{t}_{z}$ we have $\operatorname{\mathsf{Tr}}^{\infty}(\mathsf{t}_{x}\mathsf{t}_{y}^{*})=\sum_{d\in\mathcal{D}}n_{d}\gamma_{x,y^{-1},d}.$ If $y\neq x$ then P2 gives $\operatorname{\mathsf{Tr}}^{\infty}(\mathsf{t}_{x}\mathsf{t}_{y}^{*})=0$, and if $y=x$ then by P3, P5, and P13 we have $\operatorname{\mathsf{Tr}}^{\infty}(\mathsf{t}_{x}\mathsf{t}_{y}^{*})=n_{d(x)}\gamma_{x,x^{-1},d(x)}=n_{d(x)}^{2}=1$. Thus $\operatorname{\mathsf{Tr}}^{\infty}(\mathsf{t}_{x}\mathsf{t}_{y}^{*})=\delta_{x,y}$ for all $x,y\in W$. It follows that for $A,B\in\mathcal{J}$ with $A=\sum a_{w}\mathsf{t}_{w}$ and $B=\sum b_{w}\mathsf{t}_{w}$ we have
$\operatorname{\mathsf{Tr}}^{\infty}(AB^{*})$ $=\sum_{x,y\in W}n_{d(x)}n_{d(y)}a_{x}b_{y}\operatorname{\mathsf{Tr}}^{\infty}(\mathsf{t}_{x}\mathsf{t}_{y}^{*})=\sum_{x\in W}a_{x}b_{x}=\operatorname{\mathsf{Tr}}^{\infty}(B^{*}A)$
and $\langle A,A\rangle^{\infty}=\sum_{w\in W}a_{w}^{2}$, and (1) and (2) follow.
Let $x,y,z\in W$. By P7 and the fact that $\gamma_{u,v,w}=\gamma_{v^{-1},u^{-1},w^{-1}}$ (for $u,v,w\in W$) we have $\gamma_{x^{-1},z,y^{-1}}=\gamma_{y^{-1},x^{-1},z}=\gamma_{x,y,z^{-1}}$ and so by (2) we have
$\langle\mathsf{t}_{x}\mathsf{t}_{y},\mathsf{t}_{z}\rangle^{\infty}$ $=\gamma_{x,y,z^{-1}}=\gamma_{x^{-1},z,y^{-1}}=\langle\mathsf{t}_{y},\mathsf{t}_{x}^{*}\mathsf{t}_{z}\rangle^{\infty}$
and (3) follows. ∎
Remark 5.3.
Theorem 5.2 is expressed in the setup for general weighted Hecke algebras. It is clear that the statement of this theorem also applies to extended affine Hecke algebras (the proof applies verbatim).
@@ 5.2 The Plancherel Theorem for $\tilde{\mathsf{A}}_{n}$
We now return to the case that $\widetilde{H}$ is the extended affine Hecke algebra of type $\tilde{\mathsf{A}}_{n}$. The Plancherel Theorem is a spectral decomposition of the canonical trace functional $\operatorname{\mathsf{Tr}}:\widetilde{H}\to\mathbb{Z}[\mathsf{q},\mathsf{q}^{-1}]$. We now explicitly describe this decomposition in type $\tilde{\mathsf{A}}_{n}$, following [2, 27]. Recall that $\mathsf{v}=-\mathsf{q}^{-1}$. Let
@@
The Plancherel Theorem is usually expressed as an analytic statement, and some comments are required for the translation between Theorem 5.4 and this analytic statement from [2, 27]. Specialise $\mathsf{q}\to q$ with $q>1$ a real number, and extend scalars of $\widetilde{H}$ to $\mathbb{C}$. Let $\mathbb{T}$ be the group of complex numbers of modulus $1$ and for $\lambda\vdash n+1$ let $dz_{\lambda}$ be normalised Haar measure on the group $\mathbb{T}_{\lambda}=\{(z_{1},z_{2},\ldots,z_{r(\lambda)})\in\mathbb{T}^{r(\lambda)}\mid z_{1}^{\lambda_{1}}z_{2}^{\lambda_{2}}\cdots z_{r(\lambda)}^{\lambda_{r(\lambda)}}=1\}.$ On specialising the “variables” $z_{1},\ldots,z_{r(\lambda)}$ in our representations to complex numbers of modulus $1$ with $z_{1}^{\lambda_{1}}z_{2}^{\lambda_{2}}\cdots z_{r(\lambda)}^{\lambda_{r(\lambda)}}=1$, properties of the Haar measure gives
@@
and $c^{(3)}(\zeta_{(3)})=1$. We have $|G_{(1,1,1)}|=6$ and $|G_{(2,1)}|=|G_{(3)}|=1$. For the analytic formulation of the Plancherel Theorem, note that if $\lambda=(3)$ then $\mathbb{T}_{\lambda}=\{z\in\mathbb{C}\mid z^{3}=1\}=\{1,\omega,\omega^{2}\}$ where $\omega=e^{2\pi i/3}$, and so the normalised Haar measure on $\mathbb{T}_{\lambda}$ is the discrete measure assigning mass $1/3$ to each atom, and we recover the analytic formulation of the Plancherel Theorem for $\tilde{\mathsf{A}}_{2}$ computed explicitly in [28, Theorem 4.3].
@@ 5.3 Asymptotic Plancherel Theorem
The following asymptotic Plancherel Theorem gives a spectral decomposition of the inner product $\langle\cdot,\cdot\rangle^{\infty}$ on $\mathcal{J}$ (and hence the trace functional $\operatorname{\mathsf{Tr}}^{\infty}$ on $\mathcal{J}$) from Definition 5.1, mirroring the standard Plancherel Theorem at the asymptotic level.
Theorem 5.7.
For $A,B\in\mathcal{J}$ we have
$\langle A,B\rangle^{\infty}=\sum_{\lambda\vdash n+1}\langle A,B\rangle_{\lambda}^{\infty}\quad\text{where}\quad\langle A,B\rangle_{\lambda}^{\infty}=\frac{1}{|G_{\lambda}|}\bigg{[}\chi_{\lambda}^{\infty}(AB^{*})\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})\bigg{]}_{\mathrm{ct}}.$
Moreover, for each $\lambda\vdash n+1$ the bilinear form $\langle\cdot,\cdot\rangle_{\lambda}^{\infty}$ is an inner product on the $\mathbb{Z}$-module $\mathcal{J}_{\lambda}$, and the elements $\mathsf{t}_{w}$, $w\in\Delta_{\lambda}$, form an orthonormal basis.
The proof of Theorem 5.7 is given after the following two preliminary results. Recall the definition of the specialisation $\mathrm{sp}_{0}$ from Section 4.3.
@@
We now give the proof of the asymptotic Plancherel Theorem.
Proof of Theorem 5.7.
By bilinearity it is sufficient to prove that
$\sum_{\lambda\vdash n+1}\langle\mathsf{t}_{u},\mathsf{t}_{v}\rangle_{\lambda}^{\infty}=\delta_{u,v}\quad\text{for all $u,v\in\widetilde{W}$}.$
If $u\not\sim_{LR}v$ then $\mathsf{t}_{u}\mathsf{t}_{v}^{*}=0$ and so $\chi_{\lambda}^{\infty}(\mathsf{t}_{u}\mathsf{t}_{v}^{*})=0$, hence $\langle\mathsf{t}_{u},\mathsf{t}_{v}\rangle_{\lambda}^{\infty}=0$ for all $\lambda\vdash n+1$. Thus suppose that $u,v\in\Delta_{\mu}$ for some $\mu\vdash n+1$. Since $\langle T_{u},T_{v}\rangle=\delta_{u,v}$ the Plancherel Theorem (Theorem 5.4) gives
$\delta_{u,v}=\sum_{\lambda\vdash n+1}\langle T_{u},T_{v}\rangle_{\lambda},$
and by Corollary 5.9 each specialisation $\mathrm{sp}_{0}\big{(}\langle T_{u},T_{v}\rangle_{\lambda}\big{)}$ exists, and is zero unless $\lambda=\mu$. Thus $\delta_{u,v}=\mathrm{sp}_{0}\big{(}\langle T_{u},T_{v}\rangle_{\mu}\big{)}.$ Now, by Theorem 4.7 we have
$\chi_{\mu}(T_{u}T_{v^{-1}})=\mathsf{q}^{2\ell(\mathsf{w}_{\mu^{\prime}})}\operatorname{\mathsf{tr}}(\mathfrak{c}_{\mu}(u)\mathfrak{c}_{\mu}(v)^{*})+(\textrm{terms of strictly lower degree}),$
@@
$C_{\mu}(\mathsf{q})\frac{\chi_{\mu}(T_{u}T_{v^{-1}})}{c^{\mu}(\zeta_{\mu})c^{\mu}(\zeta_{\mu}^{-1})}$ $=\chi_{\mu}^{\infty}(\mathsf{t}_{u}\mathsf{t}_{v}^{*})\prod_{\alpha\in\Phi_{G_{\mu}}}(1-\zeta_{\mu}^{\alpha})+(\textrm{terms of degree $<0$}).$
@@
$\delta_{u,v}=\mathrm{sp}_{0}\big{(}\langle T_{u},T_{v}\rangle_{\mu}\big{)}=\frac{1}{|G_{\mu}|}\bigg{[}\chi_{\mu}^{\infty}(\mathsf{t}_{u}\mathsf{t}_{v}^{*})\prod_{\alpha\in\Phi_{G_{\mu}}}(1-\zeta_{\mu}^{\alpha})\bigg{]}_{\mathrm{ct}}=\langle\mathsf{t}_{u},\mathsf{t}_{v}\rangle_{\mu}^{\infty}$ (5.2)
as required. ∎
@@
By Corollary 4.8, it only remains to show that if $w\in\Delta_{\lambda}$ then $\deg\pi_{\lambda}(T_{w};\mathsf{B}_{\lambda})=\ell(\mathsf{w}_{\lambda^{\prime}})$. If $w\in\Delta_{\lambda}$ then by the asymptotic Plancherel Theorem (Theorem 5.7) we have $1=\langle\mathsf{t}_{w},\mathsf{t}_{w}\rangle_{\lambda}^{\infty}$ and thus $\deg\langle T_{w},T_{w}\rangle_{\lambda}=0$ (see (5.2)), and the result follows from Corollary 5.9. ∎
@@
By Proposition 4.12 the map $\psi:\mathcal{J}_{\lambda}\to\mathfrak{C}_{\lambda},\mathsf{t}_{w}\mapsto\mathfrak{c}_{\lambda}(w)$ is a surjective ring homomorphism. To prove that $\psi$ is injective, if $a=\sum_{w\in\Delta_{\lambda}}a_{w}\mathsf{t}_{w}\in\mathcal{J}_{\lambda}$ with $\psi(a)=0$ then for each $v\in\Delta_{\lambda}$ we have $\sum_{w\in\Delta_{\lambda}}a_{w}\mathfrak{c}_{\lambda}(w)\mathfrak{c}_{\lambda}(v)^{*}=0.$ Taking traces, multiplying by $\prod_{\alpha\in\Phi_{G_{\lambda}}}(1-\zeta_{\lambda}^{\alpha})$, and applying Theorem 5.7 gives $a_{v}=0$. ∎
@@
for $A,B\in\mathcal{J}$, where $d\mu_{\lambda}^{\infty}(z_{\lambda})$ is the asymptotic Plancherel measure.
Remark 5.13.
The results of this subsection verify [14, Conjecture 9.8] and [13, Conjecture 5.15] for affine type $\mathsf{A}$.